OpenStudy (anonymous):

What is the solution to the equation 3^1-2x = 2?

sam (.sam.):

\[3^{1-2 x}=2\]

sam (.sam.):

\[\begin{array}{l} 3^{1-2 x}=2 \\ \text{take the log to the base }3\text{ of both sides and rearrange it} \\ 1-2 x=\frac{\log (2)}{\log (3)} \\ \text{Subtract }1\text{ from both sides} \\ -2 x=\frac{\log (2)-\log (3)}{\log (3)} \\ \text{divide both sides by }-2 \\ x=-\frac{\log (2)-\log (3)}{2 \log (3)} \\\end{array}\]

OpenStudy (anonymous):

The options are: -0.292 -0.185 0.185 0.292

sam (.sam.):

do you have a calculator?

OpenStudy (anonymous):

no.

sam (.sam.):

you need it to calculate decimal solutions. 0.185

OpenStudy (anonymous):

Thanks!

sam (.sam.):

welcome